### v3

# Multimodal LLM With Hierarchical Mixture-of-Experts for VQA on 3D Brain MRI

Abstract.
Multiparametric 3D brain MRI (mpMRI) is central to neuroradiology, but producing tumor location, appearance, size, and involvement of critical structures for neurosurgical planning remains challenging. We introduce mpLLM, a multimodal LLM for visual question answering (VQA) on mpMRI that produces clinically interpretable tumor descriptors (e.g., volume, morphology, extent, and coarse localization) as an adjunct to clinical expertise for referring neurosurgeons. mpLLM uses a prompt-conditioned hierarchical mixture-of-experts (MoE) to fuse multiple 3D sequences via routing over modality- and token-level projection experts, enabling data-efficient end-to-end training without large-scale image–report pretraining. To address limited paired image–text supervision, we propose a synthetic VQA protocol that derives clinically grounded questions and answers from expert segmentation annotations and is validated with radiologist collaboration. Across multiple mpMRI datasets, mpLLM improves over strong medical VLM baselines by +5.5 points on average (+9.1% relative) and increases radiologist-rated clinical acceptability by +15.9 points (+46.6% relative). Our study features three main contributions: (1) the first VQA dataset for 3D brain mpMRI, (2) a hierarchical MoE architecture for joint reasoning over interrelated 3D sequences, and (3) expert-supported evidence of clinical utility. Source code is available at https://github.com/arvindmvepa/mpllm, and we will release the dataset upon publication.
Multimodal LLM, Mixture-of-Experts, VQA, Synthetic Data, Medical Imaging
copyright: noneccs: Applied computing Health informaticsccs: Computing methodologies Neural networksccs: Computing methodologies Natural language generationccs: Computing methodologies Computer vision
Figure 1. High-level comparison between LLaVA-Med and mpLLM. While LLaVA-Med uses a standard projection layer, our method uses a hierarchical MoE block which ingests both the prompt and imaging to produce prompt-conditioned vision tokens that leverage all the 3D modalities.
High-level comparison between LLaVA-Med and mpLLM. While LLaVA-Med uses a standard projection layer, our method uses a hierarchical MoE block which ingests both the prompt and imaging to produce prompt-conditioned vision tokens that leverage all the 3D modalities.
1. Introduction
Multiparametric MRI (mpMRI) plays a significant role in diagnosing, grading, treating, and assessing treatment responses for brain tumors and other intracranial lesions (Sawlani et al., 2020; Wang et al., 2022a; Cherubini et al., 2016). Producing clear, consistent descriptions of tumor size, morphology, extent, and localization is time-consuming, yet critical for downstream decision-making. As a result, a growing body of work has developed recognition and localization models to support clinical interpretation (Ghadimi et al., 2025; Rathore et al., 2018; Wang et al., 2022a; Li et al., 2023c; Osman, 2019).
A key unmet need is interactive, queryable assistance: neurosurgical planning and counseling depend not only on detecting tumors but on communicating clinically interpretable descriptors—where the lesion is, what it looks like, how large it is, and their involvement of critical structures. However, existing models have limited clinical utility because clinicians cannot effectively pose natural language queries about mpMRI. While 3D vision-language models (VLMs) have been developed for other imaging domains, current architectures do not naturally leverage the interdependencies among mpMRI modalities (Li et al., 2023a; Wu et al., 2023; Bai et al., 2024; Xin et al., 2025). Additionally, the standard multi-image approach multiplies the number of vision tokens by the number of images, which significantly increases computational constraints (Wu et al., 2023).
We introduce mpLLM, a prompt-conditioned hierarchical mixture-of-experts (MoE) for VQA over mpMRI. Our approach is an extension of the LLaVa architecture (Liu et al., 2023b). mpLLM replaces the standard single projection with a prompt-conditioned MoE projection that routes across modality-level and token-level projection experts to fuse multiple 3D sequences into an efficient representation for the language model. Unlike approaches that require training separate modality-specific encoders, we use lightweight projection functions that train end-to-end with the language model during fine-tuning.
To address limited paired image–text supervision in neuroimaging, we pair mpLLM with a synthetic VQA protocol that derives clinically grounded question–answer pairs from expert segmentation annotations, developed in collaboration with radiologists to ensure clinical relevance. In contrast to prior work, we fine-tune directly on the resulting VQA supervision via next-token prediction without pretraining on paired image–report data, and we train an end-to-end multi-task head to improve task proficiency and support reliable evaluation. In summary, our research makes these key contributions:
  1. (1)
    In collaboration with radiologists, we introduce a synthetic VQA protocol that produces the first VQA dataset for 3D brain mpMRI.
  2. (2)
    We design mpLLM, a multimodal LLM with prompt-conditioned hierarchical MoE routing that leverages interdependencies among 3D mpMRI modalities.
  3. (3)
    We provide empirical evidence of clinical utility through expert-supported validation and clinically relevant downstream evaluation.
Figure 2. Detailed overview of our mpLLM pipeline.
Detailed overview of our mpLLM pipeline.
2. Related Work
Medical vision-language models
Most vision-based medical multimodal LLMs can be broadly classified into CLIP-based discriminative models (Radford et al., 2021; Wang et al., 2022b; Eslami et al., 2023; Zhang et al., 2023a; Xu et al., 2024; Zhou et al., 2024; Huang et al., 2023) and LLM decoder-based generative models (Zhang et al., 2023b; Li et al., 2023a; Moor et al., 2023). Although discriminative models have proven helpful for various image recognition tasks, they possess limited utility in generation tasks such as VQA or report generation. Several popular generative models including MedVInt (Zhang et al., 2023b), LLaVA-Med (Li et al., 2023a), and MedFlamingo (Moor et al., 2023) share very similar architectures. However, these architectures and many others (Liu et al., 2024c; Lin et al., 2023; Li et al., 2023b; Zhu et al., 2024a; Lin et al., 2025; Zhang et al., 2025b; Nath et al., 2024; Guo et al., 2025) are designed specifically for 2D medical imaging and are not tailored to handle multiple 3D medical image modalities.
Although several 3D VLMs exist for natural images (Zhu et al., 2024b; Li et al., 2024b; Zhu et al., 2023), they require access to extremely large annotated datasets, which are often unavailable in medical contexts. While a few 3D VLMs have been developed for medical imaging, these methods have certain limitations. In one recent paper, researchers adapted the LLaVA-Med architecture to utilize spatial pooling and pretrain a 3D vision encoder with 700k radiology images (Bai et al., 2024). In another paper, researchers pretrain segmentation modules to generate brain imaging reports (Lei et al., 2024). In recent work, researchers exploit vision-language pretraining for CT report generation (Liu et al., 2023a; Chen and Hong, 2024; Blankemeier et al., 2024; Xin et al., 2025; Cao et al., 2025; Li et al., 2025a). However, these prior works assume a large paired imaging-report pretraining dataset, which is infeasible to collect and imposes a significant training burden.
Furthermore, previous methods focus on report generation instead of VQA, leading to less precise feedback regarding model strengths and weaknesses. Additionally, some models train directly on segmentation annotations (Lei et al., 2024; Rui et al., 2024), which are impractical to obtain, especially for novel use cases. Moreover, none of the previously discussed methods are tailored to handle multiple interdependent 3D image modalities, like in mpMRI, as input.
Mixture-of-experts
Previous work in MoE has concentrated on training and inference efficiency (Shazeer et al., 2017; Lepikhin et al., 2020; Fedus et al., 2022; Zoph et al., 2022; Liu et al., 2024a), transfer learning (Li et al., 2022; Zhong et al., 2022), class imbalance (Han et al., 2024), and multi-domain information (Zhang et al., 2024). There have also been earlier efforts with multimodal LLMs, covering sparsity learning (Lin et al., 2024), task interference (Shen et al., 2025), and embedding models (Li and Zhou, 2024). MoE also has various applications in the medical field. These applications include addressing missing modalities (Yun et al., 2025; Novosad et al., 2024; Liu et al., 2024d), fairness (Wang et al., 2025), pediatric care (Huy et al., 2025), parameter reduction and efficiency (Jiang et al., 2024; Nathani et al., 2024), super resolution (Lin et al., 2021), and segmentation of multimodal medical imaging (Zhang et al., 2025a; Jiang and Shen, 2024).
Related to our work, several studies have employed MoE with VLMs to select between vision encoders and vision-language projections (Li et al., 2025b; Zong et al., 2024; Wang et al., 2023; Ma et al., 2025). In CuMo (Li et al., 2025b), the authors incorporate MoE in the MLP layers in the vision encoder as well as projection function; however, they do not utilize prompt-conditioning, do not focus on combining multiple inputs, and require a complex training scheme (our approach requires a single stage of instruction fine-tuning). In MoVA (Zong et al., 2024), the authors utilize MoE with task-specific vision encoders. In practice, training vision encoders is challenging and resource intensive in the medical space (Rui et al., 2024). Improving the LLaVa projection function with a MoE block uses far less parameters and can be trained end-to-end with the LLM using next-token prediction. The previous approach also does not focus on combining multiple inputs. Finally, in 3D-MoE (Ma et al., 2025), the authors utilize MoE within the LLM and do not focus on combining multiple inputs. Additionally, no existing research has explored using MoE to project multiple interrelated 3D image modalities into the language embedding space.
Medical VQA Datasets
One of the primary challenges in report generation is evaluation: lexical metrics such as BLEU, ROUGE‑L, and BERTScore have been shown to correlate poorly with radiologist evaluations (Yu et al., 2023). In contrast, VQA allows for more granular and interpretable model evaluation. While there are several medical VQA datasets, many focus on 2D imaging (Liu et al., 2024b, 2021; He et al., 2020; Lau et al., 2018). In a prior work, researchers used a scene graph generator to generate surgical VQA (Yuan et al., 2024). In a recent work, researchers extracted multi-task questions from structured lung cancer screening data (Niu et al., 2025). While there are existing VQA datasets for 2D brain MRI (Safari et al., 2025; Shehzad et al., 2025; Peng et al., 2025), there are no existing VQA datasets for 3D brain mpMRI due to a significant lack of source data for VQA extraction. In our work, we remedy this by collaborating with radiologists to generate clinically relevant synthetic data from expert annotated segmentation annotations and leverage ChatGPT-4o for dataset augmentation. Other prior works have also leveraged commercial LLMs to generate VQA data from expert annotations and study metadata in the medical setting (Zhang et al., 2023b; Gautam et al., 2024; Butsanets et al., 2025). To contextualize dataset reliability and demonstrate practical value, we report data quality metrics, radiologist acceptability ratings of model-generated responses, and performance on clinically relevant downstream tasks.
3. Methodology
Brain mpMRI has several 3D imaging modalities. For a given modality $m$, let $I_{m}\in\mathbb{R}^{C\times D\times H\times W}$ denote the corresponding 3D volume, where $C$, $D$, $H$, and $W$ represent the channel, depth, height, and width, respectively. A 3D vision encoder $h_{\text{vis}}$ maps each modality to a sequence of image tokens $v_{m}=h_{\text{vis}}(I_{m})\in\mathbb{R}^{T\times d_{V}}$, where $T$ is the number of image tokens and $d_{V}$ is the vision embedding dimension. We apply spatial pooling to reduce the token length (reusing the symbol $T$ for the pooled length for simplicity) and then concatenate the pooled tokens across the $M$ image modalities before passing them to the hierarchical MoE.
Let $v\in\mathbb{R}^{M\times T\times d_{V}}$ denote the concatenated image modality embeddings. The hierarchical MoE projects these embeddings into the LLM space, $e=\mathrm{MoE}(v,t)\in\mathbb{R}^{M\times T\times d_{T}}$, where $\mathrm{MoE}$ denotes the hierarchical MoE block, $t$ represents the text prompt, and $d_{T}$ is the LLM embedding dimension. In practice we flatten the modality and token dimensions and provide $(t,e)$ as a soft prompt to the LLM for multi-task prediction and text generation. A detailed visualization of our approach can be seen in Figure 2.
3.1. Hierarchical mixture-of-experts for multiparametric MRI projection
In what follows, we use the term expert exclusively for projection modules that map vision features into the LLM embedding space, and the term router for the MLPs that output mixing weights over experts. Each high-level expert block therefore consists of a router together with its associated projection experts.
High-level router
Our hierarchical MoE architecture includes a high-level router $r^{(h)}$ that assigns weights over a set of high-level expert blocks $\{\mathcal{E}^{(h)}_{1},\dots,\mathcal{E}^{(h)}_{H}\}$, where $H$ is the number of high-level experts. These expert blocks operate at the image modality and image token levels. The router is implemented as a two-layer MLP that takes as input the final hidden state of the language model corresponding to the text prompt $t$. It produces a normalized weight distribution over expert blocks: $\pi^{(h)}(t)=\mathrm{softmax}(r^{(h)}(t))\in\mathbb{R}^{H}$. Since task information is embedded within the text prompt, the router implicitly infers the task, enabling high-level experts to specialize in different task proficiencies.
High-level image modality-level and image token-level experts
Our hierarchical MoE includes high-level experts operating at different granularity levels: image modality-level and image token-level. Each high-level expert block consists of a two-layer MLP low-level router $r^{(l)}$ and an associated set of low-level projection experts $\{W^{(l)}_{1},\dots,W^{(l)}_{L}\}$, where $L$ is the number of low-level experts within the block.
The image modality-level expert block uses a low-level router that takes as input the concatenated [CLS] tokens from all image modalities (e.g., T1, T2) and outputs modality-level weights over the corresponding low-level experts: $\pi^{(l)}_{\text{mod}}(v)=\sigma(r^{(l)}_{\text{mod}}(v))\in\mathbb{R}^{L}$. In contrast, the token-level expert block uses a low-level router that receives, for each token position $i$, the $i$-th image tokens from all modalities and outputs token-level weights $\pi^{(l)}_{\text{tok}}(v)=\sigma(r^{(l)}_{\text{tok}}(v))\in\mathbb{R}^{L\times T}$. As discussed in prior work (Li et al., 2024a), providing weights at different granularities improves task performance by enhancing domain generalizability.
Low-level image modality-specific and image modality-agnostic (shared) experts
Each low-level expert $W$ represents a projection transformation from the vision encoder embedding space to the LLM embedding space: $W:R^{N_{I}\times d_{I}}\rightarrow R^{N_{I}\times d_{T}}$. We utilized a simple linear transformation for the projection transformation as in the original LLaVa paper (Liu et al., 2023b). Each image modality embedding is processed through a modality-specific expert and a modality-agnostic (shared) expert. The modality-specific expert emphasizes extracting image modality-specific features, whereas the modality-agnostic expert focuses on deriving common features from all image modalities. The parameters for the modality-specific expert are unique to each image modality (T1Gd, T1, T2, and FLAIR), while those for the modality-agnostic expert are consistent across all image modalities.
Each image modality is passed through both low-level experts and then summed embedding dimension-wise. The overall formulation for the hierarchical MoE is as follows:
(1) $\mathrm{MoE}(v,t)$ $=\sum_{h=1}^{H}\pi^{(h)}_{h}(t)\,\cdot$
$\quad\sum_{m=1}^{M}\Big(\alpha^{(h)}_{m}\,W^{(h)}_{m}(v_{m})+\beta^{(h)}_{m}\,W^{(h)}_{\text{shared}}(v_{m})\Big)$
where $h$ indexes the high-level expert blocks, $m$ indexes the image modalities, and $\alpha^{(h)}_{m}$ and $\beta^{(h)}_{m}$ are the modality-specific and modality-agnostic weights produced by the corresponding low-level router (which sum to one within an expert block $h$). Our validation experiments on the GLI dataset found that the optimal number of high-level experts was 16, corresponding to the number of labels times the number of tasks.
The fused image token embeddings are combined with the text prompt token embeddings. Then, these embeddings are input into the LLM decoder at the token embedding layer for multi-task prediction and text generation.
Table 1. Comparison of task performance for all models on all datasets with accuracy metric with standard deviation.
[…]
3.2. Training objectives
3.2.1. Synthetic VQA protocol
Because of the lack of brain mpMRI VQA data, we propose a novel method of synthetic VQA generation that leverages the expert annotated brain mpMRI segmentation data. We perform the following steps: 1) compute continuous quantities from masks, 2) map to quantities to report-style discrete labels, and 3) insert discrete labels into paraphrased QA templates.
To generate relevant VQA data, we consult with radiologists to identify important topics that can be extracted from the label masks, focusing on mask volume relative to brain volume (Kaifi, 2023), brain region localization (Lau et al., 2018), shape (Ismail et al., 2018), and spread (Islam et al., 2019). For each label mask, we compute the quantities using standard formulas and validate the thresholds with synthetic masks and a subset of data. To emulate the subjectivity found in medical reports, we categorize each of the quantities based on their magnitude using terminology similar to that found in medical reports. We employ a rules-based method to assign medical terms to the quantities. We assign “N/A” if the label is not found.
Volume
To calculate the relative mask volume, we determine the number of mask pixels and divide by the number of brain pixels in the volume (which are the nonzero pixels in the skull-stripped T1 image modality). The subjective labels we use are $1\%$”, $1-5\%$”, $5-10\%$”, $10-25\%$”, $25-50\%$”, and “507550-75%”.
Region
For the BraTS GLI volumes, we use the Nibabel python library (Abraham et al., 2014) to register the volumes to the AAL atlas (version SPM12) (Rolls et al., 2020) and extract the following brain regions: “frontal”, “parietal”, “occipital”, “temporal”, “limbic”, “insula”, “subcortical”, and “cerebellum”. For the BraTS MET and GoAT volumes, we register the volumes to the LPBA40 atlas (in SRI24 space) (Shattuck et al., 2008) and extract the following brain regions: “frontal”, “parietal”, “occipital”, “temporal”, “limbic”, “insula”, “subcortical”, “cerebellum”, and “brainstem”. The percent coverages of the segmentation masks with the atlases are 67.3%, 70.9%, and 57.7% for the GLI, MET, and GoAT datasets respectively.
Shape
We first quantify each mask’s overall size and compute classical 3‑D shape metrics (sphericity, elongation, flatness, solidity, compactness). If the mask is tiny, it is classified as “focus”; otherwise, we classify it as “round,” “oval,” “elongated,” or “irregular” by comparing its sphericity and elongation values to empirically chosen thresholds that correspond to near‑sphere, mildly flattened, and strongly stretched geometries.
Spread
We identify all disconnected islands, noting the largest as the “core,” and compute what proportion of the total mask volume it occupies. If there is only one island, the pattern is “single lesion”; if multiple islands are present but the core retains $\geq$85% of the volume, it is described as “core with satellite lesions”; otherwise, when no dominant island exists, the distribution is marked “scattered lesions.”
Question-answer pair generation
After computing the previous quantities for each label mask, we create a dataset that simulates the natural variability of human input. First, we consider all combinations of the four major tasks to create multi-task question-answer pairs. After we have the 15 question-answer pair types, we use ChatGPT-4o to generate approximately 3000 perturbations of each question-answer pair (without affecting the label and answer term) that emulates the language a radiologist would use. We also add question-answer pairs with partially out-of-scope and completely out-of-scope tasks to improve the model’s self-awareness of its capabilities. Thus, for each label and mpMRI in each dataset, we sample four multitask question-answer pairs without replacement such that each major task is addressed in at least one question-answer pair, one partially out-of-scope question-answer pair, and one completely out-of-scope question-answer pair. Examples of generated question-answer pairs can be seen in the Appendix in Table 11. The answers are used as supervision for next-token prediction for the multimodal LLM.
3.2.2. Multi-task heads
For increased task proficiency and more accurate task evaluation, we train a multi-task head end-to-end with the multimodal LLM. After providing the soft-prompt to the multimodal LLM, we extract the hidden state from the last layer and apply task-specific heads (which consist of a single linear layer) to generate multi-task predictions. For volume, shape, spread, and out-of-scope task identification, the task is multi-class classification, and the associated loss is categorical cross-entropy; whereas for region localization, the task is multi-label classification, and the associated loss is multi-label binary cross-entropy. These losses are added to the next-token prediction loss to produce our multi-task loss:
(2) $\mathcal{L}$ $=\mathcal{L}_{\textrm{Next-token}}+\mathcal{L}_{\textrm{Volume}}+\mathcal{L}_{\textrm{Region}}$
$\quad+\mathcal{L}_{\textrm{Shape}}+\mathcal{L}_{\textrm{Spread}}+\mathcal{L}_{\textrm{Out-of-scope}}$
4. Experiments
4.1. Datasets details
For our synthetic VQA protocol, we leverage the 2024 Brain Tumor Segmentation (BraTS) challenge (LaBella et al., 2024), which provides a standardized benchmarking environment for automated brain tumor segmentation. All datasets comprise of co-registered multiparametric MRI scans (T1, T1Gd, T2, FLAIR) at 1mm3 resolution, skull-stripped and manually annotated by experts. To enable fair comparison and manage GPU memory, all BraTS sequences were resampled to $32\times 256\times 256$. This allowed for compatibility with baseline methods, such as M3D (Bai et al., 2024) and Med3DVLM (Xin et al., 2025). We consider three challenges in BraTS: GLI, MET and GoAT. The challenges are collected from over ten institutions and encompass diverse pathological contexts and imaging protocols.
GLI (Adult Glioma Post Treatment)
Dataset focuses on post-treatment diffuse glioma segmentation and consists of multi-institutional routine post-treatment clinically-acquired multiparametric mpMRI scans of glioma. The task requires the delineation of enhancing tumor (ET), non-enhancing tumor core (NETC), surrounding FLAIR hyperintensity (SNFH), and resection cavity (RC) (de Verdier et al., 2024).
MET (Brain Metastases)
Dataset contains a retrospective compilation of treatment-naive brain metastases mpMRI scans obtained from various institutions under standard clinical conditions. The challenge addresses the segmentation of small metastatic lesions using a 3-label system (NETC, SNFH, ET) and demonstrates variable tumor component distribution across cases (Moawad et al., 2024).
GoAT (Generalizability Across Tumors)
Dataset assesses algorithmic generalizability across different tumor types (i.e., different number of lesions per scan, lesion sizes, and locations in the brain), institutions (i.e., different MRI scanners, acquisition protocols), and demographics (i.e., different age, sex, etc.). The challenge uses consistent labels (necrosis, edema/invaded tissue, and enhancing tumor) despite varying tumor morphology to evaluate algorithm adaptability to new disease types with limited training data (de Verdier et al., 2024; Moawad et al., 2024; LaBella et al., 2023; Kazerooni et al., 2024; Adewole et al., 2023).
To generate the train, validation, and test sets, we randomly sample 80%, 10%, and 10% from the imaging studies. For GLI we generated 31,104, 4,176, and 3,624 question-answer pairs for the train, validation, and test sets based on 1,621 mpMRIs. For MET, we generated 9,090, 1,368, and 1,260 question-answer pairs for the train, validation, and test sets, based on 651 mpMRIs. For GoAT, we generated 19,440, 2,430, and 2,448 question-answer pairs for the train, validation, and test sets, based on 1351 mpMRIs. All the generated questions pass a rules-based filter which checks for proper formatting. Additionally, all the generated test set questions pass an automated data cleaning step using ChatGPT-4o to ensure high quality evaluation.
Clinical validation
We collaborated with two radiologists who annotated 20 mpMRIs from the BraTS-GLI test set, 10 mpMRIs from the BraTS-MET test set, and 10 mpMRIs from the BraTS-GoAT test set with questions spanning four tasks and four findings for the BraTS-GLI dataset and three findings for the BraTS-MET and BraTS-GoAT datasets, yielding 560 questions. We used 10 annotated mpMRIs from BraTS-GLI for validation to improve the task label thresholds for the synthetic data. We used the other 30 annotated mpMRIs to evaluate the inter-annotator agreement between the synthetic data and each radiologist, the model predictions and each radiologist, and both radiologists (Table 15 and 16).
To assess the quality of our synthetic questions, a radiologist evaluated the clarity of 160 synthetic questions from 10 mpMRIs from the BraTS-GLI test set using a binary scoring system (1 = valid, 0 = invalid). Prior to the LLM automated data cleaning, the synthetic questions achieved a 92.2% validity rate. After automated data cleaning, the synthetic questions achieved a 95.9% validity rate. More details about the datasets can be found in Appendix A and  B.
4.2. Experimental settings
Models
In our experiments, we utilized the Phi-3-Mini-4K-Instruct LLM. We also explored utilizing the Llama models and chose Phi-3 because of the increased efficiency and negligible performance benefits of the Llama models. For versatility and generality, we utilize the 3D Vision Transformer (3D ViT) (Dosovitskiy et al., 2020) as the vision encoder and use medically pretrained weights (Bai et al., 2024).
Training
We fine-tune the multimodal LLM using the loss defined in Equation 2 on the VQA training dataset. We freeze the vision encoder while unfreezing the hierarchical MoE and LLM weights. We train the model on the train dataset for 2 epochs. The LLM is trained with LoRA, setting $r$ to 16 and alphaalpha to 32, with a dropout of 0.1. We employ a cosine learning rate scheduler that starts at a learning rate of $2.0\times 10^{-4}$.
Baseline models
We compare our approach to several baseline models, including LLaVA-Med (Li et al., 2023a), M3D (Bai et al., 2024), Merlin (Blankemeier et al., 2024), Med3DVLM (Xin et al., 2025), and RadFM (Wu et al., 2023). To process the multiple 3D MRI image modalities, we use a multi-image approach, in which we concatenate the image tokens generated from each MRI image modality from a shared projection layer and vision encoder (Wu et al., 2023). Because LLaVA-Med is not implemented with a 3D vision encoder, to ensure a fair comparison, we test it with our model’s vision encoder (Bai et al., 2024). Similar to our method, the vision encoder is frozen and only the projection layer and LLM are trainable. To provide a comparison to our model’s multi-task classifier, which is trained end-to-end with the rest of our framework, we independently train a new multi-task classifier. We use a Phi3 language model with multi-task heads to predict the multi-task outputs given the prompt and text generation. The model is trained on our train dataset and had 99.8% accuracy on the validation set. Other hyperparameter settings mirror our method as closely as possible to ensure a fair comparison.
Evaluation
For evaluating the models’ task proficiency, we use accuracy for volume, shape, spread, and out-of-scope tasks, and per-label accuracy for the region task. We estimate the standard deviation using 500 bootstrap resamples.
Computing environment
All our experiments were mainly conducted using a single NVIDIA A100 GPU on an internal cluster. Training our model on the GLI dataset took approximately 10 hours.
Table 2. Ablation study on the MoE architecture on the GLI validation set with accuracy metric.
[…]
rowsep=0.95pt Modality-level MoE Token-level MoE Prompt-based MoE weights Task Mean 63.3 64.1 64.4 65.5
Table 3. Comparison of MoE-based and shared expert approach based on task on the GLI validation set with accuracy metric
Method Volume Region Shape Spread
Shared expert 53.2 82.8 59.7 57.5
MoE-based 57.6 83.5 60.9 59.9
Table 4. Comparison of MoE-based and shared expert approach based on finding on the GLI validation set with accuracy metric
Method ET SNFH NETC RC
Shared expert 63.2 70.8 71.3 48.7
MoE-based 65.1 72.5 75.1 49.9
Table 5. Comparison of MoE approaches based on task on the GLI validation set with accuracy metric
Method Volume Region Shape Spread
Modality-level MoE 53.7 83.4 59.7 59.4
Token-level MoE 56.0 83.5 60.0 58.1
Prompt-based MoE 57.6 83.5 60.9 59.9
Table 6. Comparison of MoE approaches based on finding on the GLI validation set with accuracy metric
Method ET SNFH NETC RC
Modality-level MoE 63.5 71.1 73.9 48.6
Token-level MoE 64.0 72.2 72.2 49.9
Prompt-based MoE 65.1 72.5 75.1 49.9
Table 7. Comparison of task means with and without region information on the GLI validation set with accuracy metric.
Model Task Mean without Region Scores
Without Region Information 59.5
With Region Information 64.1
Table 8. Radiologist acceptance rate comparison between mpLLM and M3D.
Model Radiologist Acceptance Rate (%)
M3D 34.1
mpLLM 50.0
Table 9. Comparison of model performance for differentiating primary gliomas versus secondary metastatic lesions
Model Accuracy AUROC
M3D 88.5 95.5
mpLLM 95.6 99.0
4.3. Results
All model results across the evaluated datasets are presented in Table 1. Our model consistently achieves strong performance across all task categories and datasets, outperforming the second-best model by an average margin of 5.5 points (+9.1% relative to the strongest baseline). We note a small drop in performance relative to other models on the GoAT dataset. This dataset is extremely challenging because it evaluates generalizability based on several factors with limited training data. In our experiments, we also noticed that the top three models had above a 99.8% accuracy on out-of-scope task identification, which suggests our dataset was effective at hallucination mitigation.
4.4. Ablation studies
An ablation study on the MoE architecture is in Table 2. Image modality-level and token-level high-level MoE experts perform better than the single projection layer baseline approach. A prompt-conditioned weighted combination of the different high-level experts performs the best.
Fine-grained results comparing our MoE-based approach and a single shared expert are in Table 3 and Table 4. The more complex multimodal reasoning is helpful for all tasks and findings, and especially helpful for the volume task (+4.4 points and 8.3% relative to the baseline) and NETC finding (+3.8 points and 5.3% relative to the baseline). Fine-grained results comparing modality-level MoE, token-level MoE, as well prompt-conditioned MoE on tasks as well as findings are in Table 5 and Table 6. Token-level MoE is stronger in the volume, region, and shape tasks while modality-level MoE is stronger in the spread task. For findings, token-level MoE is stronger with the ET, SNFH, and RC findings while modality-level MoE is stronger with the NETC finding. The prompt-conditioned MoE excels in all of them, suggesting that based on the question, the model is able to accurately combine the optimal token-level MoE or modality-level MoE blocks.
4.5. Clinical perspective
From a neuroradiology standpoint, an important use case for mpLLM is as an adjunct to clinical expertise: producing standardized, interpretable tumor descriptors (e.g., size, morphology, extent, and coarse localization) for referring neurosurgeons. This framework also provides a pathway to more fine-grained future outputs, such as involvement with specific gyri or eloquent cortex. A unique technical challenge in neuro-oncology imaging is robustness to atypical anatomy and our results suggest the proposed approach remains effective across highly variable lesion appearances.
To quantify clinical utility under this framing, we evaluate (i) agreement between model predictions and radiologist annotations, (ii) radiologist acceptability of generated responses, and (iii) clinically relevant downstream performance. Agreement results are shown in Table 16. We observe that mpLLM is comparable to radiologist-level consistency on the volume and shape tasks, indicating strong reliability for two core descriptors commonly communicated in clinical workflows. Performance is lower on region and spread, which we attribute primarily to higher noise in the synthetic labeling pipeline for these attributes and to the inherent difficulty of localization/extent estimation when anatomy is atypical or deformed. This result highlights a practical constraint for real-world deployment: model performance is bounded by label quality and by the stability of the underlying definitions.
We further probe the role of localization by appending region information to the prompt (Table 7), yielding an approximately 5% increase on other task scores. This suggests (i) localization provides useful signal for estimating morphology/extent and (ii) even relatively noisy region labels can improve other clinically relevant descriptors, warranting their inclusion. This aligns with clinical practice, where coarse localization is often communicated alongside morphology/extent.
Because many descriptors are subjective, we collaborated with a radiologist to assess whether generated answers are clinically reasonable. We created a clinical validation set of 208 questions from 13 GLI test cases, focusing on volume, region, shape, and spread (questions 1–4 in Appendix A). We compared mpLLM against M3D by asking the radiologist to label each response as sufficient or lacking. mpLLM achieves a 50.0% acceptability rating and improves clinical acceptability by +15.9 points (+46.6% relative; Table 8). These findings support the role of mpLLM as a reporting adjunct for tumor descriptors.
Finally, to emphasize downstream usefulness, we consider a clinically important task: differentiating primary gliomas from secondary metastatic lesions using imaging-derived descriptors (volume, region, shape, spread). Because our models are trained with guard rails and indicate if tasks are outside of their scope, we fit a logistic regression classifier on model-generated features from the combined training set of BraTS GLI and MET. On the test set (Table 9), mpLLM improves over M3D by +7 accuracy points and +4 AUROC points, suggesting that the learned structured descriptors capture discriminative imaging patterns relevant to neuro-oncology workflows.
5. Conclusion
We present mpLLM, a multimodal LLM for mpMRI VQA that uses prompt-conditioned hierarchical MoE routing over modality- and token-level projection experts, enabling efficient end-to-end fine-tuning without paired image–report pretraining. We pair this model with a synthetic VQA pipeline derived from expert segmentation annotations, developed in collaboration with a radiologist, to produce standardized tumor descriptors (e.g., volume, morphology, extent, and coarse localization) that can serve as an adjunct to clinical expertise for referring neurosurgeons. Across BraTS GLI/MET/GoAT, mpLLM improves over strong medical VLM baselines by an average of +5.5 points and increases radiologist-rated clinical acceptability by +15.9 points, with ablations validating the contributions of modality/token experts and prompt-conditioned routing. These findings suggest that structured, interpretable outputs can remain reliable across heterogeneous lesion appearances, while also highlighting the importance of label quality for more challenging localization tasks. Future work includes extending to open-ended VQA and report generation, evaluating robustness in settings with atypical anatomy, conducting broader multi-reader validation, and performing fairness analyses.
6. Limitations and Ethical Considerations
This work uses publicly available, fully de-identified BraTS datasets, minimizing risks to patient privacy and data security. Our synthetic VQA are generated from segmentation annotations, and both the generated questions and model outputs underwent radiologist review to mitigate typical risks of synthetic supervision. Nonetheless, fairness and bias remain open concerns: synthetic prompts and limited demographic metadata can yield models that underperform for underrepresented groups or clinical scenarios. The model is intended for research only and must not be used for autonomous clinical decision-making; it is designed to abstain on out-of-scope queries, and any deployment would require prospective, multi-site validation under qualified clinical oversight. In future work, we will expand evaluations to demographically diverse cohorts where available, document dataset composition and known limitations, and incorporate explicit fairness analyses and bias-mitigation strategies alongside robustness and calibration assessments.
7. GenAI Disclosure
We used large language models (LLMs) to (i) improve the clarity and style of the manuscript, (ii) brainstorm refinements to the MoE-based architecture and dataset-construction procedures, (iii) draft code prototypes for selected ideas, and (iv) find potentially relevant related work. All LLM outputs were reviewed and verified by the authors before inclusion.
Appendix A Additional information regarding dataset
[…]
rowsep=0.95pt Dataset # questions # mpMRI # unique questions # unique answers GLI 38,904 1,621 38,023 36,773 MET 11,718 651 11,607 11,284 GoAT 24,318 1,351 23,859 23,223
Table 10. Statistics for the synthetic VQA datasets.
In the following, we will describe the formulas used to derive the shape and spread descriptors for our synthetic VQA protocol. Let $M\subset\mathbb{Z}^{3}$ be a binary mask of lesion voxels sampled with spacing $\mathbf{s}=(s_{x},s_{y},s_{z})\,[\mathrm{mm}]$ (typically $s_{x}=s_{y}=s_{z}=1$). Write $\Delta V=s_{x}s_{y}s_{z}$ for the physical volume of one voxel and $\lvert M\rvert$ for the number of lesion voxels.
Total volume
$V_{\text{tot}}=\lvert M\rvert\,\Delta V\;[\mathrm{mm}^{3}].$
Multiplicity
We define 26-connectivity on $M$ (scipy ‘ndimage.label‘ with a unit “ball” structuring element) and record $N_{\!c}$ and $M_{1},\dots,M_{N_{\!c}}$.
Spread
Let the core component index be $i^{\star}=\arg\max_{i}V_{i}$. Define
$f_{\text{core}}\;=\;\frac{V_{i^{\star}}}{V_{\text{tot}}}\in[0,1].$
$\text{spread}=\begin{cases}\text{``single lesion''}&N_{\!c}=1,\\[2.0pt] \text{``core with satellite lesions''}&N_{\!c}>1,\;f_{\text{core}}\geq 0.85,\\[2.0pt] \text{``scattered lesions''}&\text{otherwise.}\end{cases}$
For each component $M_{i}$:
Component surface area
Marching cubes (scikit-image ‘measure.marching_cubes‘) produces a triangular mesh $(\mathcal{V}_{i},\mathcal{F}_{i})$ in real-world coordinates. The mesh area (which we describe as the surface area) is
$A_{i}=\!\!\!\sum_{(p,q,r)\in\mathcal{F}_{i}}\!\!\!\tfrac{1}{2}\bigl\|(q-p)\times(r-p)\bigr\|_{2}.$
Component volume
$V_{i}=\lvert M_{i}\rvert\,\Delta V$.
Component sphericity
$\Phi_{i}\;=\;\frac{\pi^{1/3}\,(6V_{i})^{2/3}}{A_{i}}.$
Component compactness
$C_{i}\;=\;\frac{A_{i}}{V_{i}}.$
Component principal-axis statistics
Assemble voxel coordinates $\mathbf{x}_{j}=(x_{j},y_{j},z_{j})\in\mathbb{R}^{3}$ for $j\in M_{i}$. The covariance matrix $\Sigma_{i}=\frac{1}{\lvert M_{i}\rvert}\sum_{j}(\mathbf{x}_{j}-\bar{\mathbf{x}})(\mathbf{x}_{j}-\bar{\mathbf{x}})^{\!\top}$ yields eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}>0$.
Component elongation
$E_{i}\;=\;\sqrt{\lambda_{1}/\lambda_{2}}.$
Component flatness
$F_{i}\;=\;\sqrt{\lambda_{3}/\lambda_{2}}.$
Component solidity
A convex hull (scipy ‘ConvexHull‘) provides volume $V^{\mathrm{hull}}_{i}$;
$S_{i}\;=\;\frac{V_{i}}{V^{\mathrm{hull}}_{i}}.$
Metric aggregation
$(\Phi,E,F,S,C)=\begin{cases}(\Phi_{i^{\star}},E_{i^{\star}},F_{i^{\star}},S_{i^{\star}},C_{i^{\star}})&N_{\!c}=1\text{ or }f_{\text{core}}\geq 0.7,\\[4.0pt] \displaystyle\frac{1}{N_{\!c}}\sum_{i=1}^{N_{\!c}}(\Phi_{i},E_{i},F_{i},S_{i},C_{i})&\text{otherwise.}\end{cases}$
Shape
Convert the continuous metrics to one of five categories:
$\text{shape}=\begin{cases}\text{``focus''}&V_{\text{tot}}<0.1\;\text{cm}^{3}\quad(\;V_{\text{tot}}\times 10^{-3}<0.1\;)\\[2.0pt] \text{``round''}&\Phi\geq 0.80\;\land\;E\leq 1.4,\\[2.0pt] \text{``oval''}&0.50\leq\Phi<0.80\;\land\;1.4<E\leq 3.0,\\[2.0pt] \text{``elongated''}&E>3.0,\\[2.0pt] \text{``irregular''}&\text{otherwise.}\end{cases}$
The thresholds were set empirically on 10 annotated mpMRIs from BraTS-GLI for validation set and match radiologists’ qualitative intuition of near-spherical, mildly flattened, and strongly stretched geometries. All computations are implemented in Python using scipy, scikit-image, numpy, and ndimage as shown in the listing above.
Question augmentation details
We use ChatGPT to generate question augmentations of our multitask dataset. For generating question augmentations for the standard multi-task prompts, we first provide this prompt “Please produce hundred alternative wordings that a radiologist may use for the following question and answer. Please include everything surrounded by curly braces {} as they are because they are placeholders. Please generate the reworded question starting with “Q:” and reworded answer starting with “A:” and separate each generated question-answer pair with a newline. Please do not produce any additional text.” and append this to each of the multitask prompts below. We produce 40 repetitions with a temperature of $1.0$, top p of $1$, and model “gpt-4o-mini-2024-07-18”.
  1. (1)
    Q: How large is the volume covered by {label}? A: The overall volume of {label} is {volume}.
  2. (2)
    Q: Which region(s) of the brain is {label} located in? A: The {label} is located in {regions}.
  3. (3)
    Q: What is the shape of {label}? A: The shape of {label} is {shape}.
  4. (4)
    Q: How spread out is {label}? A: The spread of {label} is {spread}.
  5. (5)
    Q: How large is the volume of {label} and where is it located? A: The overall volume of {label} is {volume}, and it is located in {regions}.
  6. (6)
    Q: How large is the volume of {label} and what is its shape? A: The overall volume of {label} is {volume}, and its shape is described as {shape}.
  7. (7)
    Q: How large is the volume of {label} and how spread out is it? A: The overall volume of {label} is {volume}, and it is characterized as {spread}.
  8. (8)
    Q: In which region is {label} and what is its shape? A: The {label} is located in {regions}, and its shape is described as {shape}.
  9. (9)
    Q: In which region is {label} and how spread out is it? A: The {label} is located in {regions}, and it is characterized as {spread}.
  10. (10)
    Q: What is the shape of {label} and how spread out is it? A: The shape of {label} is described as {shape}, and it is characterized as {spread}.
  11. (11)
    Q: What is the volume, region, and shape of {label}? A: The overall volume of {label} is {volume}, it is located in {regions}, and its shape is described as {shape}.
  12. (12)
    Q: What is the volume, region, and spread of {label}? A: The overall volume of {label} is {volume}, it is located in {regions}, and it is characterized as {spread}.
  13. (13)
    Q: What is the volume, shape, and spread of {label}? A: The overall volume of {label} is {volume}, its shape is described as {shape}, and it is characterized as {spread}.
  14. (14)
    Q: What is the region, shape, and spread of {label}? A: The {label} is located in {regions}, its shape is described as {shape}, and it is characterized as {spread}.
  15. (15)
    Q: What is the volume, region, shape, and spread of {label}? A: The overall volume of {label} is {volume}, it is located in {regions}, its shape is described as {shape}, and it is characterized as {spread}.
For generating question augmentations for the partially out-of-scope multi-task prompts, we first provide this prompt “Please produce hundred alternative wordings that a radiologist may use for the following question and answer and incorporate an additional clinical task or tasks which the model cannot solve in the reworded question. These can be before, after, or interspersed between the other tasks (please make sure to vary the order and number of out-of-scope tasks). Do not mention that the model cannot answer these in the question; however, indicate that the model cannot answer that part of the question in the reworded answer (potentially using different phrasings). The model can describe the volume, brain region, shape, and spread of {label} which is the region of interest. Please include everything surrounded by curly braces {} as they are because they are placeholders. Please generate the reworded question starting with “Q:” and reworded answer starting with “A:” and do not produce any additional text.” and append this to each of the multitask prompts above. We produce 10 repetitions with a temperature of $1.0$, top p of $1$, and model “gpt-4o-2024-08-06”.
For generating question augmentations for completely out-of-scope prompts, we first provide this prompt “Please produce a hundred questions (with one or more tasks) that a radiologist may ask that the model does not have information to answer. The model can describe the volume, brain region, shape, and spread of {label} which is the region of interest. Please include {label} in the question but do not include anything else with curly braces. In the answer, please indicate the model cannot answer the question (potentially using different phrasings). Please generate the question starting with “Q:” and answer starting with “A:” and do not produce any additional text.” We produce 10 repetitions with a temperature of $1.0$, top p of $1$, and model “gpt-4o-mini-2024-07-18”.
After generating the question augmentations, we apply a rules-based quality check: questions and answers are properly split with ”Q:” and ”A:” respectively, contents within the curly braces are retained for easy formatting with Python, and that the responses are in English. Then, for each finding and mpMRI in each dataset, we sample four multitask questions without replacement such that each major task is addressed in at least one question, one partially out-of-scope question, and one completely out-of-scope question. Examples of generated question types can be seen in Table 11. Additionally, all the generated test set question candidates pass an automated data cleaning step using ChatGPT (model “gpt-4o-2024-08-06”) to ensure high quality evaluation. See the prompts below for the three different question types.
  1. (1)
    Determine whether the reworded question is a faithful, clinically appropriate rephrasing of the original question. Please ignore valid as this is a placeholder. Output VALID or INVALID with no additional text.
  2. (2)
    Determine whether the reworded question is a faithful, clinically appropriate rephrasing of the original question. The reworded question also has an additional query not in the original question. Please also determine whether the added query is a valid medical question. Please ignore label as this is a placeholder. Output VALID or INVALID with no additional text. Example 1: Original Q: How large is the volume of label and what is its shape? Reworded Q: What is the mass and form of label, and engage in patient counseling? INVALID Example 2: Original Q: How large is the volume of label and what is its shape? Reworded Q: Can you estimate the volume and layout of label, and administer a flu vaccine? VALID
  3. (3)
    Determine whether the reworded question is a valid medical question. Please ignore valid as this is a placeholder. Output VALID or INVALID with no additional text.
To assess the quality of our synthetic questions, a radiologist evaluated the clarity of 160 synthetic questions from 10 mpMRIs from the BraTS-GLI test set using a binary scoring system (1 = valid, 0 = invalid). Prior to the LLM automated data cleaning, the synthetic questions achieved a 92.2% validity rate. After automated data cleaning, the synthetic questions achieved a 95.9% validity rate.
After the application of our synthetic VQA protocol, the percentage frequency of each task per question for all the generated datasets can be seen in Table 12, Table 13, and Table 14.
Table 11. Qualitative examples illustrating the four question types generated by our pipeline.
Type Question Answer
Single‑task Could you identify the shape of the Resection Cavity? The shape of the Resection Cavity is irregular.
Multi‑task How do you quantify the volume of Surrounding Non‑enhancing FLAIR hyperintensity, and what does its distribution look like? The total volume of Surrounding Non-enhancing FLAIR hyperintensity is described as 1–5%, characterized as core with satellite lesions.
Partial‑out-of-scope What areas does Surrounding Non‑enhancing FLAIR hyperintensity encompass, what is its volume, and synthesize findings into a comprehensive care plan? Surrounding Non-enhancing FLAIR hyperintensity has a volume of 1–5% in cerebellum, frontal and parietal. Synthesizing this into a care plan is outside my domain.
Out-of-scope How do genetic factors influence the development of Non‑Enhancing Tumor in adolescents? I cannot provide information regarding the influence of genetic factors on the development of Non-Enhancing Tumor in adolescents.
Table 12. Percentage frequency of each task label per question for the GLI dataset
[…]
rowsep=0.95pt Task Label name Label frequency Volume Unspecified 52.4 N/A 12.3 <<1% 12.7 1-5% 13.1 5-10% 5.2 10-25% 3.8 25-50% 0.5 50-75% 0.0 Region Unspecified 53.2 N/A 12.2 frontal 23.8 parietal 20.8 occipital 13.2 temporal 17.2 limbic 21.7 insula 14.7 subcortical 14.9 cerebellum 2.9 Shape Unspecified 52.4 N/A 12.3 focus 1.0 round 4.7 oval 6.9 elongated 0.4 irregular 22.4 Spread Unspecified 53.4 N/A 12.2 single lesion 6.9 core with satellite lesions 20.9 scattered lesions 6.5 Out-of-scope Not out-of-scope 66.7 Out-of-scope 33.3
Table 13. Percentage frequency of each task label per question for the MET dataset
[…]
rowsep=0.95pt Task Label name Label frequency Volume Unspecified 51.8 N/A 8.5 <<1% 24.7 1-5% 9.2 5-10% 2.7 10-25% 2.7 25-50% 0.4 50-75% 0.0 Region Unspecified 53.0 N/A 17.1 frontal 19.5 parietal 16.1 occipital 14.5 temporal 14.5 limbic 9.2 insula 6.3 subcortical 7.3 cerebellum 12.8 brainstem 4.1 Shape Unspecified 51.4 N/A 8.3 focus 2.8 round 13.5 oval 4.3 elongated 0.2 irregular 19.5 Spread Unspecified 53.1 N/A 7.9 single lesion 7.4 core with satellite lesions 12.9 scattered lesions 18.7 Out-of-scope Not out-of-scope 66.7 Out-of-scope 33.3
Table 14. Percentage frequency of each task label per question for the GoAT dataset
[…]
rowsep=0.95pt Task Label name Label frequency Volume Unspecified 52.2 N/A 2.2 <<1% 10.3 1-5% 18.6 5-10% 8.9 10-25% 7.2 25-50% 0.6 Region Unspecified 52.9 N/A 2.3 frontal 30.0 parietal 23.4 occipital 17.5 temporal 29.7 limbic 29.8 insula 25.9 subcortical 27.9 cerebellum 9.8 brainstem 8.0 Shape Unspecified 51.9 N/A 2.1 focus 0.5 round 6.1 oval 3.2 elongated 0.1 irregular 36.0 Spread Unspecified 53.3 N/A 1.9 single lesion 5.3 core with satellite lesions 32.7 scattered lesions 6.7 Out-of-scope Not out-of-scope 66.7 Out-of-scope 33.3
Appendix B Additional Ablation results
Additional ablation results validating the number of high-level experts, softmax versus sigmoid for summing lower-level experts, and concatenation versus element-wise summing of vision tokens are in Table 17, Table 18, and Table 19 respectively. In Table 20, we see a comparison of our model performance trained with our multi-task loss versus the next-token prediction baseline loss. There is a significant performance improvement with our multi-task loss.
To qualitatively evaluate our architecture, we construct all 60 template task prompts from our GLI dataset (four findings $\times$ 15 task combinations = 60 template prompts) and input them into our model’s high-level router to generate high-level expert weight vectors. We then calculate the correlation between these weight vectors and generate a heatmap, which is in Figure 3. There’s high correlation between expert weight vectors within the same finding, suggesting similar image features are extracted. For findings that are closer anatomically, such as non-enhancing tumor core and enhancing tissue, there is also relatively high correlation between the expert weight vectors, which suggests similar extracted image features. For findings like resection cavity and surrounding FLAIR hyperintensity that are more diverse anatomically from the other findings, there’s much lower correlation, which suggests more dissimilar extracted features.
Appendix C Reproducibility statement
We use a publicly available dataset and detail the full data-generation pipeline in Sections 3.2.1 and 4.1, with additional information in Appendix A. We provide the github link our model code in the abstract, and report all experimental settings and computational resources in Section 4.2.
Table 15. Inter-annotator agreement (accuracy) between synthetic data and radiologists
[…]
rowsep=0.95pt Agreement between synthetic data and Volume Region Shape Spread Radiologist 1 51.0 74.2 56.0 39.0 Radiologist 2 49.0 75.7 43.0 48.0
Table 16. Inter-annotator agreement (accuracy) between model predictions and radiologists and between both radiologists
[…]
rowsep=0.95pt Agreement between Volume Region Shape Spread Model and Radiologist 1 43.0 70.9 47.0 23.0 Model and Radiologist 2 43.0 71.2 49.0 38.0 Both Radiologists 38.0 85.5 50.0 62.0
Table 17. Model performance comparison with different number of high-level experts on the GLI validation set with accuracy metric.
[…]
rowsep=0.95pt, colsep=3.5pt Number of blocks Task Mean 12 64.5 16 65.5 20 65.0
Table 18. Model performance with softmax versus sigmoid for summing lower-level experts on the GLI validation set with accuracy metric.
[…]
rowsep=0.95pt, colsep=3.5pt Method Task Mean sigmoid 64.8 softmax 65.5
Table 19. Comparison of projection and fusion methods on the GLI validation set with accuracy metric.
Projection Method Fusion Method Task Mean
Shared expert concatenation 52.4
Shared expert sum 51.1
Shared expert learned weighted sum 63.3
MoE-based sum 65.5
Table 20. Model comparison with multi-task loss on the GLI validation set with accuracy metric.
[…]
rowsep=0.95pt, colsep=3.5pt Method Task Mean mpLLM without multi-task loss 59.5 mpLLM with multi-task loss 65.5
Figure 3. Heatmap for correlation between high-level expert weight vectors for standard prompts in the GLI dataset. NETC = non-enhancing tumor core, ET = enhancing tissue, SNFH = surrounding FLAIR hyperintensity, RC = resection cavity.
Heatmap for correlation between high-level expert weight vectors for standard prompts in the GLI dataset. NETC = non-enhancing tumor core, ET = enhancing tissue, SNFH = surrounding FLAIR hyperintensity, RC = resection cavity.